# Tracking the Unstable: Appearance-Guided Motion Modeling for Robust Multi-Object Tracking in UAV-Captured Videos

### v1

@@ IV-C5 Robustness to Large Displacements
IV-C6 Generality on Other Trackers
We apply our design to the representative JDE-based trackers, including UAVMOT [27] and STCMOT [28]. As presented in Table VI, all of these trackers benefit from the integration of AMC and MTC. For example, STCMOT achieves a notable gain of +7.0% IDF1 and +2.1% MOTA on VisDrone2019. Similar performance improvements are observed for UAVMOT, with IDF1 and MOTA increasing by +4.4% and +2.4%, respectively. As shown in Fig. VI, trackers enhanced with the proposed components consistently outperform their original counterparts under varying frame intervals. Such improvements across diverse trackers and scenarios further demonstrate the generality and effectiveness of our approach.
IV-D Visualization
Frequent viewpoint changes in UAV-captured videos often induce substantial variations in object appearance and position, posing significant challenges to multi-object trackers. As illustrated in Fig. 7, we evaluate the tracking performance of our method under these adverse conditions. Both the baseline model and MM-tracker [8] suffer from frequent identity switches and tracking failures, reflecting limited capability in coping with dynamic viewpoint changes. In contrast, the proposed AMOT demonstrates robust performance and maintains track identity consistency under such challenging conditions.
V Conclusion

### v2

@@ Introduction
Table 6: Results of applying AMC and MTC components to various JDE-based methods.
Generality on Other Trackers.
We apply our design to the representative JDE-based trackers, including UAVMOT (Liu et al. 2022) and STCMOT (Ma et al. 2024). As presented in Table 6, all of these trackers benefit from the integration of AMC and MTC. For example, STCMOT achieves a notable gain of +7.0% IDF1 and +2.1% MOTA on VisDrone2019. Similar performance improvements are observed for UAVMOT, with IDF1 and MOTA increasing by +4.4% and +2.4%, respectively. Such improvements across diverse trackers and scenarios further demonstrate the generality and effectiveness of our approach.
Visualization
Frequent viewpoint changes in UAV-captured videos often induce substantial variations in object appearance and position, posing significant challenges to multi-object trackers. As illustrated in Figure 6, we evaluate the tracking performance of our method under these adverse conditions. Both the baseline model and MM-tracker (Yao et al. 2025) suffer from frequent identity switches and tracking failures, reflecting limited capability in coping with dynamic viewpoint changes. In contrast, the proposed AMOT demonstrates robust performance and maintains track identity consistency under such challenging conditions.
Conclusion
@@
Acknowledgments
We gratefully acknowledge the support of Prof. Haorui Zuo. The work was partially supported by Frontier Research Fund of the Institute of Optics and Electronics, Chinese Academy of Sciences (C24K003).
Appendix A -Supplementary Material-
In this work, we propose AMOT, an advanced multi-object tracker built upon the joint detection and embedding (JDE) architecture (Zhang et al. 2021), augmented with two novel components: the appearance-motion consistency (AMC) matrix and the motion-aware track continuation (MTC) module. AMOT is specifically designed to enhance the robustness and accuracy of identity assignment under challenging UAV-captured videos. Comprehensive experiments and qualitative evaluations on UAV benchmarks are included in the supplementary material.
Details of Hyperparameter Settings
We conduct detailed ablation studies to analyze the impact of the hyperparameter settings used in the AMC matrix and the MTC module.
Scale Factor $\sigma$ in the AMC Matrix.
Here, we construct the AMC matrix $\mathbf{C}_{AMC}$ using a Gaussian kernel that integrates both forward and backward spatial distances, defined as:
where $\mathbf{D}_{f}$ and $\mathbf{D}_{b}$ represent the forward and backward spatial distances, respectively, $\sigma$ is a scale factor that controls the spatial sensitivity. We employ a Gaussian kernel to normalize the bi-directional spatial distance into the range [0,1], thereby constructing a smooth and bounded affinity matrix. As illustrated in Figure 7, a smaller $\sigma$ (e.g., $\sigma=1$ or $\sigma=2$) results in a sharper decay of $\mathbf{C}_{AMC}$ with increasing spatial distance, reflecting higher sensitivity to spatial proximity. While a larger $\sigma$ (e.g., $\sigma=8$ or $\sigma=9$) leads to a smoother decay, indicating reduced sensitivity. Different values of $\sigma$ reflect varying sensitivities to spatial proximity.
To further examine the impact of $\sigma$, we analyze the association performance under different settings, as reported in Table 7. The IDF score exhibits an increasing trend with larger $\sigma$ values at first, but deteriorates beyond a certain point due to reduced sensitivity to spatial proximity. The optimal balance is achieved when $\sigma=5$, which yields the highest IDF and a competitive number of identity switches (IDs). Thus, we adopt $\sigma=5$ as the default setting in all experiments.
[…]
1 74.1 55.1 793 142 78727 73949 338
2 74.2 55.1 793 142 78821 73930 311
3 74.4 55.1 793 142 78925 73900 276
4 74.6 55.1 794 142 78931 73907 260
5 74.7 55.1 794 142 78928 73968 272
6 74.7 55.1 793 142 78723 74151 288
7 74.5 55.1 792 143 78346 74256 329
8 74.0 55.1 793 144 78161 74353 356
9 73.9 55.2 792 144 77619 74622 381
10 73.5 55.3 789 146 77052 74883 401
Table 7: Impact of the scale factor $\sigma$ on the association performance.
Figure 7: Impact of $\sigma$ on the sensitivity of $\mathbf{C}_{AMC}$ as the distance increases.
Predefined Threshold $\lambda$ in MTC Module.
As shown in Figure 8, we propose the MTC module to determine whether to reactivate unmatched tracks by computing the Euclidean distance $d$ between the appearance-guided predicted center $\mathbf{c}_{reid}$ with the Kalman-based predicted center $\mathbf{c}_{kf}$. This distance is formulated as:
$d$ $=\left\|\mathbf{c}_{reid}-\mathbf{c}_{kf}\right\|_{2}.$ (13)
Then, we compare $d$ with a predefined threshold $\lambda$ to determine the activation of unmatched tracks. In Table 8, we present the impact of different predefined threshold $\lambda$. Although increasing $\lambda$ leads to a reduction in false negatives (FN), it also introduces more false positives (FP). Nevertheless, the best overall tracking performance is achieved at $\lambda=3$, where both MOTA and IDF1 reach their highest values, indicating an effective balance among FP, FN, and IDs. Consequently, we adopt $\lambda=3$ as the default setting in all experiments.
[…]
1 74.5 55.1 784 143 77651 75132 291
2 74.6 55.1 790 142 78469 74402 277
3 74.7 55.1 794 142 78928 73968 272
4 74.7 55.0 795 142 79266 73807 270
6 74.7 54.9 795 142 79660 73676 268
8 74.6 54.9 795 142 79830 73645 268
10 74.6 54.9 795 142 79911 73616 269
15 74.5 54.8 795 142 80126 73574 271
20 74.5 54.8 796 142 80268 73539 271
40 74.4 54.5 798 142 81440 73388 271
Table 8: Impact of the predefined threshold $\lambda$ on the tracking performance.
Figure 8: Overview of the MTC module. At frame $t$, track T#2 could be associated with detection D#1, but track T#1 remains unmatched due to a missed detection. Subsequently, we determine whether to reactivate T#1 by comparing its appearance-guided predicted center $\mathbf{c}_{reid}$ with the Kalman-based predicted center $\mathbf{c}_{kf}$.
Cost matrix VisDrone2019 UAVDT
[…]
$\mathbf{C}_{uni}^{\scriptscriptstyle(1)}$ 61.4 46.0 1063 74.7 55.1 272
$\mathbf{C}_{uni}^{\scriptscriptstyle(2)}$ 56.8 45.3 2480 73.8 55.1 446
$\mathbf{C}_{uni}^{\scriptscriptstyle(3)}$ 61.2 46.1 1138 74.3 55.0 305
Table 9: Impact of different combinations of the unified cost matrix.
Figure 9: The average object displacement is calculated over all sequences and objects under different frame intervals.
Unified Cost Matrix Formulations.
Here, we explore several formulations for combining different cost matrices, including the AMC matrix $\mathbf{C}_{AMC}$, IoU matrix $\mathbf{C}_{IOU}$, and appearance similarity matrix $\mathbf{C}_{App}$, and ultimately construct a unified cost matrix $\mathbf{C}_{uni}$ for data association. $\mathbf{C}_{uni}$ aims to leverage complementary information from these individual matrices to better measure the affinity between detections and tracks. Besides $\mathbf{C}_{uni}^{\scriptscriptstyle(1)}$, we also investigate the following two formulations:
$\mathbf{C}_{uni}^{(2)}=1-(1-\mathbf{C}_{IOU})\cdot(1-\mathbf{C}_{AMC}\cdot\mathbf{C}_{App}),$ (14)
$\mathbf{C}_{uni}^{(3)}=1-(1-\mathbf{C}_{AMC}\cdot\mathbf{C}_{IOU})\cdot(1-\mathbf{C}_{AMC}\cdot\mathbf{C}_{App}).$ (15)
As shown in Table 9, $\mathbf{C}_{uni}^{\scriptscriptstyle(1)}$ consistently achieves better tracking performance, demonstrating the effectiveness of its design. While $\mathbf{C}_{uni}^{\scriptscriptstyle(2)}$ demonstrates worse performance, this may be due to the AMC matrix failing to compensate for the instability of the IoU matrix. Such results indicate that the AMC matrix effectively models the joint relationship between visual similarity and spatial coherence.
Further Evaluation on Large Displacements
In Figure 9, we present the average object displacement calculated over all sequences and objects in the VisDrone2019 and UAVDT test sets under increasing frame intervals. As the frame interval increases, the average displacement exhibits a clear upward trend, indicating that longer temporal gaps introduce greater motion uncertainty and pose challenges for multi-object trackers. The subsection titled “Robustness to Large Displacements” in the manuscript has validated the robustness of our AMOT in addressing such challenging scenarios.